### v4

@@ I Introduction
    We publish a large dataset that captures all traffic generated by 8 different application types, clearly reveal how the traffic is labeled, and share the domain names associated with all relevant and background traffic.
@@
    We address the presence of background traffic by designing a classification framework that can handle such traffic in the wild.
  3. 3.
    We investigate the trade-off between performance and coverage in traffic classification and subsequently design a classification framework with well-calibrated confidence scores that can help achieve different operating points in this trade-off based on the confidence threshold.
@@
Selenium
Web Browsing Google, Baidu, Wikipedia, Yandex, Naver, Docomo, Samsung, Weather, Microsoft, DuckDuckGo, eBay, ESPN, CNN, Etsy, AOL, USPS, Yahoo 314 104
@@
Manual
VoIP/Video Call Skype, Facebook, Zoom 129 67
@@ IV-D Comprehensive Domain Name-Based Labeling Scheme
To accommodate the presence of background traffic without adversely affecting classification performance for application-specific traffic, we introduce a new class for the former. If no substring in the domain name indicates its relevance to the corresponding application session, we label it as background. This is different from the state-of-the-art flow-based labeling scheme that only labels flows with relevant SNI and ignores the rest of the traffic, such as Orange ’20 that labels 119,565 out of 343,228 flows [1]. If we separately label the background traffic and include it as an additional class to train the model, we get the following result in Table IV.
@@ VI Gaussian Mixture Model Classification Framework
Using Gaussian Mixture Model (GMM) clustering to determine how far a sample lies from a known cluster is more reliable mean to assess the uncertainty [9]. However, in order to effectively cluster the samples, the raw time series feature vectors need to be first encoded into learned representation vector using deep learning. Prior work shows that supervised contrastive loss produces representation vectors that are better for classification than the output logits learned through cross‑entropy loss [5]. Therefore, we use the supervised contrastive loss [5] to train the deep learning encoder. To make the BiLSTM model compatible with the supervised contrastive loss, we remove the softmax output layer from the original architecture such that it outputs a 64-dimensional embedding vector. For FS-Net, the original loss is the sum of classification cross-entropy loss and reconstruction cross-entropy loss. We modify the original loss by replacing the classification loss with supervised contrastive loss. Also, we change the final layer of FS-Net to output a 64-dimensional embedding vector.

### v5

@@ I Introduction
    We address the presence of background traffic by designing a classifier that can handle such traffic in the wild.
@@
    We investigate the trade-off between performance and coverage in traffic classification and subsequently design a classification framework that can help achieve different operating points in this trade-off based on the confidence threshold.
@@
Selenium Web Browsing Google, Baidu, Wikipedia, Yandex, Naver, Docomo, Samsung, Weather, Microsoft, DuckDuckGo, eBay, ESPN, CNN, Etsy, AOL, USPS, Yahoo 314 104
@@
Manual VoIP/Video Call Skype, Facebook, Zoom 129 67
@@ IV-D Comprehensive Domain Name-Based Labeling Scheme
To accommodate the presence of background traffic without adversely affecting classification performance for application-specific traffic, we introduce a new class for the former. If no substring in the domain name indicates its relevance to the corresponding application session, we label it as background. If we separately label the background traffic and include it as an additional class to train the model, we get the following result in Table IV.
@@ VI Gaussian Mixture Model Classification Framework
Using Gaussian Mixture Model (GMM) clustering to determine how far a sample lies from a known cluster is more reliable mean to assess the epistemic uncertainty [9]. However, in order to effectively cluster the samples, the raw time series feature vectors need to be first encoded into learned representation vector using deep learning. Prior work shows that supervised contrastive loss produces representation vectors that are better for classification than the output logits learned through cross‑entropy loss [5]. Therefore, we use the supervised contrastive loss [5] to train the deep learning encoder. To make the BiLSTM model compatible with the supervised contrastive loss, we remove the softmax output layer from the original architecture such that it outputs a 64-dimensional embedding vector. For FS-Net, the original loss is the sum of classification cross-entropy loss and reconstruction cross-entropy loss. We modify the original loss by replacing the classification loss with supervised contrastive loss. Also, we change the final layer of FS-Net to output a 64-dimensional embedding vector.